OpenStudy (anonymous):

Solve 7xsqauared-20x=3

OpenStudy (anonymous):

7x^2 -20x -3 = 0 is factor able into (7x + 1) ( x-3) = 0 so by the product property of multiplication, each factor may equal zero, so you have 7x + 1 = 0 or x - 3 = 0 ... Solve each for x to get x= -1/7 or x= 3